OpenStudy (anonymous):

kindly solve this in special products: (1/2r^2-1/4r^2-r^3)^2…

OpenStudy (anonymous):

(r^4(1+16r^2))/16

OpenStudy (anonymous):

\[\left(\frac{r^2}{4}-r^3\right)^2=\frac{r^4}{16}-\frac{r^5}{2}+r^6=\frac{1}{16} r^4 (4 r-1)^2 \]